# references.bib

### v1

@@ 5.2.2. Reactive Approaches
By the continuous observation of system metrics, reactive approaches may adjust resource partitioning in reaction to changing system state.
@@
navarro-torres_balancer_2023 (navarro-torres_balancer_2023) identify changes in program phase by observing cache and memory bandwidth usage behaviors. Such changes trigger reevaluations of the current allocation scheme for LLC and memory bandwidth, respectively. The two resource dimensions are controlled by separate heuristic schemes, that work together to balance resource congestion. This controller is called Balancer, which aims to improve overall system performance and thread fairness.
@@
HyPart (park_hypart_2018) is a dynamic memory bandwidth partitioning controller with a focus on fairness proposed by park_hypart_2018 that uses workload profiling as a base for its explorative approach to resource partitioning. Profiling workload behavior allows the scheme to eliminate potential states from the state space it uses to iteratively and heuristically converge on some optimum partitioning scheme. The process is repeated should the state of the system change. By expanding upon this work, the authors later propose CoPart (park_copart_2019). In this extended work, the controller is able to optimize for performance in addition to fairness. It is also extended to partition and manage LLC allocation instead of solely managing memory bandwidth.
@@
xu_dcat_2018 (xu_dcat_2018) propose dCat, a dynamic LLC partitioning controller that aims to improve the performance of system workloads. This is done through heuristic workload phase detection based on cache access, instruction retirement, and instructions per cycle (IPC) metrics. Workloads are classified into the classes reclaim, receiver, donor, keeper, streaming, and unknown based on their current cache utilization behavior. Allocation changes are triggered by applications transitioning between states due to changes in cache access behavior.
@@
pons_cache-poll_2022 (pons_cache-poll_2022) aim to improve system fairness and performance by identifying cache-polluting processes and limiting their ability to impact cache-sensitive applications with Cache-Poll. Their approach is a heuristic algorithm that classifies applications based on system metrics into one of three classes. Cache sensitive, mildly cache sensitive, and cache insensitive. Based on the said classification, Cache-Poll then partitions available LLC in order to improve performance and fairness.
@@
By building upon their work presented in (farina_assessing_2022), farina_enabling_2023 (farina_enabling_2023) propose a heuristic for estimating and guaranteeing memory access for critical cores while not significantly sacrificing the performance of non-critical ones. Their approach is based on observation and analysis of the un-core metric read pending queue (RPQ) found on Intel processors.
@@
roy_satori_2021 roy_satori_2021 increase system throughput and fairness of multiple BE jobs competing over system resources. They do this by proposing Satori, a resource partitioning controller utilizing Bayesian optimization to optimize cache and memory bandwidth allocations based on a novel objective function using system metrics.
@@
By observing a tenant VM’s ability to significantly reduce the performance of co-located VMs, chen_alita_2020 (chen_alita_2020) argue that a dynamic resource management system is needed to improve fairness among tenants. Their proposal called Alita identifies resource polluters and limits their availability to resources that are under pressure. The identification of such polluters is done solely through the use of hardware and OS-level metrics.
@@
xiang_dcaps_2018 (xiang_dcaps_2018) approach the problem of cache partitioning by searching the large space of possible partitioning combinations for a near-optimal configuration. They do this by proposing an algorithm based on the simulated annealing optimization scheme that optimizes workload performance. This scheme is called dynamic allocation with partial sharing (DCAPS), and operates in user space, using only system metrics.
@@
Intel’s Data Direct I/O (DDIO) technology enables the transfer of data directly between network controllers and LLC, completely bypassing main memory and increasing performance. However, this may introduce additional LLC congestion. In an attempt to mitigate this, yuan_dont_2021 (yuan_dont_2021) propose IAT. IAT is a heuristic cache partitioning scheme that resembles a Mealy finite state machine (FSM). Workloads are classified into states that map onto states of the FSM. Based on system metrics, workloads may change state and thus trigger LLC reallocation. IAT is DDIO aware and may allocate part of the LLC exclusively to DDIO to improve system performance.
@@
Through observation of system metrics, yi_mt2_2022 (yi_mt2_2022) aim to improve system performance on platforms with non-volatile memory (NVM) by regulating memory bandwidth with their MT2 controller. The kernel-space controller allows administrator control of its throttling groups, which are based on the Linux kernel’s cgroups feature. By continuously updating the bandwidth throttling settings based on application bandwidth usage, overall system performance may be improved.
@@
To protect against cache timing channel attacks between hardware threads, applications, or VMs, yao_leveraging_2019 proposes a system that monitors cache behavior and heuristically identifies pairs of workloads with suspiciously correlated activity (yao_leveraging_2019). Any suspected spy-victim pair are separated by strictly separating their available cache.
@@ 5.2.3. Combined Approaches
By classifying workloads based on cache usage behavior, li_pfa_2022 (li_pfa_2022) propose a performance and fairness-aware LLC partitioning scheme called PFA. Classification is performed through the observation of cache hit rates of retired CPU instructions, after which a prediction of LLC requirements is made. During operation, performance slack is monitored and partitioning is adjusted accordingly. Improvements are observed in both performance and fairness using their controller, but it is noted that further improvements should be achievable by tuning memory bandwidth in combination with LLC partitioning.
@@
With a clear focus on fairness, garcia-garcia_lfoc_2019 (garcia-garcia_lfoc_2019) propose LFOC, a controller that classifies workloads by their cache behavior. The classes defined are light sharing, streaming, and sensitive. LLC is then partitioned accordingly using a heuristic cache-clustering algorithm to increase overall system fairness. This work is later extended in LFOC+ (saez_lfoc_2022) where an improved pair-clustering strategy is used to further improve system fairness.
@@
By dividing the LLC into two partitions, pons_improving_2018 (pons_improving_2018) attempt to improve system performance and fairness. Workloads are profiled offline and classified as either cache sensitive or cache insensitive based on how they respond to changes in cache availability, after which they are assigned to the corresponding LLC partition. By dynamically “moving the goalpost”, i.e. shrinking one partition and increasing the other, they aim to decrease system congestion. Their heuristic approach is called critical-aware or simply referred to as CA. This work is later extended in CPA (pons_phase-aware_2020) which uses instructions per cycle and cache hit rate metrics to determine the phase of a workload and allocates LLC accordingly. Here, LLC partitioning is not limited to just two partitions.
@@
Through their use of deep reinforcement learning (DRL), chen_drlpart_2021 (chen_drlpart_2021) present a resource allocation controller called DRLPart that partitions the resources of a system among workloads using a reward prediction model. The goal of the allocations is to reduce resource congestion and increase workload throughput. To validate its effectiveness, a comparison with two older controllers (patel_clite_2020; xiang_dcaps_2018) is carried out.
@@
By allowing the compiler to create hints about the cache behavior of programs, chatterjee_com-cas_2023 (chatterjee_com-cas_2023) are able to make proactive, just-in-time, LLC allocation decisions. Their proposed scheme, Com-CAS, makes use of a compilation component based on LLVM and a runtime component that works in unison to achieve improved system performance.
@@
With the goal of mitigating resource under-utilization in large-scale Kubernetes clusters, zhang_zeus_2021 detail their resource management scheme Zeus (zhang_zeus_2021). Zeus intercepts Kubernetes API requests and modifies allocation requests in accordance with its observed node-level metrics. On each node, LLC is split into two partitions. One for LC workloads, and one for BE jobs. This allows the controller to minimize interference and thus increase average, cluster-wide, CPU utilization from $15\%$ to $60\%$.
@@
To improve the performance of accelerated ML workloads, zhu_kelp_2019 leverage non-uniform memory access (NUMA) node subdomains in their Kelp (zhu_kelp_2019) runtime. Each system socket is split into two subdomains, one for high-priority ML workloads and one for low-priority tasks. This separation in combination with heuristic resource partitioning reduces task interference such that ML workload throughput is increased.
@@ 5.2.4. Workload Clustering Approaches
Themis (tang_themis_2023) by tang_themis_2023 increases QoS and fairness of guest VMs by observing their memory bandwidth usage behavior. The VMs are classified into groups using the Euclidean K-Means algorithm, where the number of groups equal the number of available COS on the system. The controller then proactively partitions LLC and MB by following a heuristic algorithm. The classification and partitioning scheme is reevaluated over time by continuously observing VM bandwidth usage. The controller may also react to saturation of the memory bus by moving VMs to other NUMA nodes if available.
By observing a core’s LLC cache miss percentage and number of retired instructions, gifford_dna_2021 (gifford_dna_2021) classify workloads into clusters depending on the program phase they are deemed to currently be in. These clusters are then allocated LLC ways and memory bandwidth555 Memory bandwidth is allocated using MemGuard (yun_memguard_2013), a software-based approach. in order to maximize the performance of a system based on a heuristic algorithm. They call their approach DNA. A deadline-aware version that is able to take a real-time workload’s execution deadlines into consideration called DADNA is also presented. Their experimental evaluation of the controller only compares DNA and DADNA against (xu_holistic_2019), another controller previously proposed by the same authors.
With the goal of increasing overall system performance, xiang_emba_2019 (xiang_emba_2019) propose EMBA. EMBA is a controller that clusters workloads using a heuristic hierarchical clustering algorithm based on their usage of the memory bus. This classification is then used to heuristically and iteratively achieve a memory bandwidth throttle scheme that improves overall performance.
With system fairness as their goal, selfa_application_2017 (selfa_application_2017) propose a set of LLC allocation policies that are based on workload clustering using the K-Means algorithm (hartigan_algorithm_1979). The classification is done based on the number of stalled CPU cycles due to L2 cache misses. Then, the LLC is partitioned accordingly which results in an improvement of overall system fairness.
By clustering workloads using a support vector machine (SVM) based machine learning model, qiu_machine-learning-based_2022 are able to classify workload behavior using only system metrics. This is the first phase of their Clustering-and-Allocation (C&A) (qiu_machine-learning-based_2022) LLC partitioning scheme. Using the aforementioned workload classification as input, a Bayesian optimization-based approach is used to quickly and efficiently reach a near-optimum partitioning scheme. In their own evaluation, C&A is found to increase throughput by up to $26.2\%$ in a best-case scenario.
@@ 5.2.5. Feature Abstractions
Due to the nature of hardware features such as Intel CAT, they are limited to bare-metal deployments with direct access to hardware. In an attempt to bring LLC allocation to virtualized environments, xu_vcat_2017 (xu_vcat_2017) propose vCAT. vCAT is a system that exposes virtual cache partitioning to VMs, which is later mapped to physical LLC partitions on the host. By the extension of the functionality of hypervisors, guests are able to give hints regarding their workloads’ requirements, while the host remains in control.
@@ 5.2.6. Summary
Hardware QoS enforcement features may be used to great effect and tuned in many ways even when only system metrics are observed. Throughput may be increased by $26.2\%$ (qiu_machine-learning-based_2022) or $18.5\%$ (chen_drlpart_2021), performance increased by $7.1\%$ (navarro-torres_balancer_2023), or fairness improved by $44\%$ (pons_cache-poll_2022) compared to an unmanaged system.
Many publications classify workloads with regard to their utilization and sensitivity to a shortage of shared system resources. This categorization is then often used to dictate how resources are partitioned among workloads. Other papers try to predict how changes in resource allocation would impact the performance of workloads through machine learning and mathematical approaches. Conversely, reactive controllers have been proposed that continuously profile the workloads and reactively adjust resource allocations as requirements change over time. Additionally, publications focused on security (liu_catalyst_2016), compiler inserted cache behavior hints (chatterjee_com-cas_2023), auctioning-based partitioning (funaro_ginseng_2016), and LLC management on the data-center scale (shahrad_provisioning_2021), and more have been discussed.
@@ 5.3. Application Aware Approaches
To better understand how shared resource allocation affects the real-world performance of workloads, controllers may monitor key performance indicators (KPIs) of applications. Such metrics may include request throughput, average latencies, tail latencies, and more. Every workload’s KPIs are different and manual configuration of metric probes is therefore often required. Resource management controllers need to be aware of how to measure an application’s metrics, as well as their target values and the importance of one metric compared to another. Many workloads have service-level objectives (SLOs) to adhere to which allow application-aware resource managers to minimize the resource availability of such workloads while ensuring they still operate within their targets. Thus, freeing more resources to maximize the performance of other co-located workloads.
We group application-aware approaches into four categories: Proactive, reactive, combined, and performance slack management approaches. All identified papers that fit the application-aware category are listed in Table 5.
@@
(jeatsa_casy_2022) Cloud Performance ML Resource Management Controller for FaaS
(stewart_performance_2022) Generic Performance ML Short-term LLC Allocation Controller
(noll_accelerating_2018) DBMS Performance LLC Management for DBMS
@@
(nikas_dicer_2019) Generic Utilization Heuristic LLC Partitioning Scheme
(chintapalli_restrain_2022) VNFs Performance Heuristic VNF Resource Management
(duan_improving_2021) Generic Performance ML Classification and Heuristic Scheduling
(duan_improving_2022) Generic Performance ML Classification and Heuristic Scheduling
(patel_clite_2020) Cloud Performance Bayesian Optimization Based Algorithm
(pons_effect_2022) Generic Performance Analysis of App Sensitivity to Shared Resources
(gureya_generalizing_2021) Cloud Performance Heuristic Multi-Socket Memory Management
(papadakis_improving_2017) Generic Performance Heuristic LLC Partitioning Scheme
(qin_locpart_2017) Generic Performance Heuristic LLC Partitioning Scheme
@@
(qiu_slo_2021) Generic Performance Profiling & Feedback Based Management
(nejat_coordinated_2020) Generic Energy Efficiency Modeling of Workload Behavior
(zhao_rhythm_2020) Cloud Utilization Sub-service LLC Management Scheme
(tootoonchian_resq_2018) VNFs Utilization Greedy Profiling Based LLC Allocation
(metsch_intent-driven_2023) Cloud Performance ML and Intent-driven Resource Management
(li_rldrm_2020) VNFs Utilization Deep Reinforcement Learning LLC Management
(zhang_libra_2021) Generic Performance Heuristic Memory Bandwidth Management
(chintapalli_ravin_2023) VNFs Utilization Heuristic VNF and Shared Resource Allocation
(xu_holistic_2019) RT Performance Heuristic LLC and BW Management Algorithm
@@
(zhu_dirigent_2016) Generic Utilization Execution Time Prediction Algorithm
(chen_parties_2019) Cloud Utilization Heuristic Slack Management Algorithm
(lo_heracles_2015) Cloud Performance Heuristic Slack Management Algorithm
(nejat_cooperative_2022) Generic Energy Efficiency Heuristic Slack Management Algorithm
(pang_adaptive_2021) Cloud Energy Efficiency Multi-level Resource Management Scheme
(chen_avalon_2019) Generic Utilization Runtime Prediction and LLC Management

### v2

@@ 5.2.2. Reactive Approaches
Reactive approaches adjust resource partitioning schemes in response to changes in workload behaviors and system metrics. They do not attempt to predict future resource requirements, but rather react to changes in the system state as they occur.
@@
(navarro-torres_balancer_2023, 67) (navarro-torres_balancer_2023, 67) identify changes in program phase by observing cache and memory bandwidth usage behaviors. Such changes trigger reevaluations of the current allocation scheme for LLC and memory bandwidth, respectively. The two resource dimensions are controlled by separate heuristic schemes, that work together to balance resource congestion. This controller is called Balancer, which aims to improve overall system performance and thread fairness.
@@
HyPart (park_hypart_2018, 75) is a dynamic memory bandwidth partitioning controller with a focus on fairness proposed by (park_hypart_2018, 75) that uses workload profiling as a base for its explorative approach to resource partitioning. Profiling workload behavior allows the scheme to eliminate potential states from the state space it uses to iteratively and heuristically converge on some optimum partitioning scheme. The process is repeated should the state of the system change. By expanding upon this work, the authors later propose CoPart (park_copart_2019, 74). In this extended work which we previously highlighted, the controller is able to optimize for performance in addition to fairness. It is also extended to partition and manage LLC allocation instead of solely managing memory bandwidth.
@@
(xu_dcat_2018, 111) (xu_dcat_2018, 111) propose dCat, a dynamic LLC partitioning controller that aims to improve the performance of system workloads. This is done through heuristic workload phase detection based on cache access, instruction retirement, and instructions per cycle (IPC) metrics. Workloads are classified into the classes reclaim, receiver, donor, keeper, streaming, and unknown based on their current cache utilization behavior. Allocation changes are triggered by applications transitioning between states due to changes in cache access behavior.
@@
(pons_cache-poll_2022, 78) (pons_cache-poll_2022, 78) aim to improve system fairness and performance by identifying cache-polluting processes and limiting their ability to impact cache-sensitive applications with Cache-Poll. Their approach is a heuristic algorithm that classifies applications based on system metrics into one of three classes. Cache sensitive, mildly cache sensitive, and cache insensitive. Based on the said classification, Cache-Poll then partitions available LLC in order to improve performance and fairness.
@@
By building upon their work presented in (farina_assessing_2022, 27), (farina_enabling_2023, 28) (farina_enabling_2023, 28) propose a heuristic for estimating and guaranteeing memory access for critical cores while not significantly sacrificing the performance of non-critical ones. Their approach is based on observation and analysis of the un-core metric read pending queue (RPQ) found on Intel processors.
@@
(roy_satori_2021, 89) (roy_satori_2021, 89) increase system throughput and fairness of multiple BE jobs competing over system resources. They do this by proposing Satori, a resource partitioning controller utilizing Bayesian optimization to optimize cache and memory bandwidth allocations based on a novel objective function using system metrics.
@@
By observing a tenant VM’s ability to significantly reduce the performance of co-located VMs, (chen_alita_2020, 15) (chen_alita_2020, 15) argue that a dynamic resource management system is needed to improve fairness among tenants. Their proposal called Alita identifies resource polluters and limits their availability to resources that are under pressure. The identification of such polluters is done solely through the use of hardware and OS-level metrics.
@@
(xiang_dcaps_2018, 108) (xiang_dcaps_2018, 108) approach the problem of cache partitioning by searching the large space of possible partitioning combinations for a near-optimal configuration. They do this by proposing an algorithm based on the simulated annealing optimization scheme that optimizes workload performance. This scheme is called dynamic allocation with partial sharing (DCAPS), and operates in user space, using only system metrics.
@@
Intel’s Data Direct I/O (DDIO) technology enables the transfer of data directly between network controllers and LLC, completely bypassing main memory and increasing performance. However, this may introduce additional LLC congestion. In an attempt to mitigate this, (yuan_dont_2021, 117) (yuan_dont_2021, 117) propose IAT. IAT is a heuristic cache partitioning scheme that resembles a Mealy finite state machine (FSM). Workloads are classified into states that map onto states of the FSM. Based on system metrics, workloads may change state and thus trigger LLC reallocation. IAT is DDIO aware and may allocate part of the LLC exclusively to DDIO to improve system performance.
@@
Through observation of system metrics, (yi_mt2_2022, 115) (yi_mt2_2022, 115) aim to improve system performance on platforms with non-volatile memory (NVM) by regulating memory bandwidth with their MT2 controller. The kernel-space controller allows administrator control of its throttling groups, which are based on the Linux kernel’s cgroups feature. By continuously updating the bandwidth throttling settings based on application bandwidth usage, overall system performance may be improved.
@@
To protect against cache timing channel attacks between hardware threads, applications, or VMs, (yao_leveraging_2019, 114) proposes a system that monitors cache behavior and heuristically identifies pairs of workloads with suspiciously correlated activity (yao_leveraging_2019, 114). Any suspected spy-victim pair are separated by strictly separating their available cache.
@@ 5.2.3. Combined Approaches
By classifying workloads based on cache usage behavior, (li_pfa_2022, 58) (li_pfa_2022, 58) propose a performance and fairness-aware LLC partitioning scheme called PFA. Classification is performed through the observation of cache hit rates of retired CPU instructions, after which a prediction of LLC requirements is made. During operation, performance slack is monitored and partitioning is adjusted accordingly. Improvements are observed in both performance and fairness using their controller, but it is noted that further improvements should be achievable by tuning memory bandwidth in combination with LLC partitioning.
@@
With a clear focus on fairness, (garcia-garcia_lfoc_2019, 33) (garcia-garcia_lfoc_2019, 33) propose LFOC, a controller that classifies workloads by their cache behavior. The classes defined are light sharing, streaming, and sensitive. LLC is then partitioned accordingly using a heuristic cache-clustering algorithm to increase overall system fairness. This work is later extended in LFOC+ (saez_lfoc_2022, 90) where an improved pair-clustering strategy is used to further improve system fairness.
@@
By dividing the LLC into two partitions, (pons_improving_2018, 80) (pons_improving_2018, 80) attempt to improve system performance and fairness. Workloads are profiled offline and classified as either cache sensitive or cache insensitive based on how they respond to changes in cache availability, after which they are assigned to the corresponding LLC partition. By dynamically “moving the goalpost”, i.e. shrinking one partition and increasing the other, they aim to decrease system congestion. Their heuristic approach is called critical-aware or simply referred to as CA. This work is later extended in CPA (pons_phase-aware_2020, 79) which uses instructions per cycle and cache hit rate metrics to determine the phase of a workload and allocates LLC accordingly. Here, LLC partitioning is not limited to just two partitions.
@@
Through their use of deep reinforcement learning (DRL), (chen_drlpart_2021, 16) (chen_drlpart_2021, 16) present a resource allocation controller called DRLPart that partitions the resources of a system among workloads using a reward prediction model. The goal of the allocations is to reduce resource congestion and increase workload throughput. To validate its effectiveness, a comparison with two older controllers (patel_clite_2020, 76, 108) is carried out.
@@
By allowing the compiler to create hints about the cache behavior of programs, (chatterjee_com-cas_2023, 12) (chatterjee_com-cas_2023, 12) are able to make proactive, just-in-time, LLC allocation decisions. Their proposed scheme, Com-CAS, makes use of a compilation component based on LLVM and a runtime component that works in unison to achieve improved system performance.
@@
With the goal of mitigating resource under-utilization in large-scale Kubernetes clusters, (zhang_zeus_2021, 121) detail their resource management scheme Zeus (zhang_zeus_2021, 121). Zeus intercepts Kubernetes API requests and modifies allocation requests in accordance with its observed node-level metrics. On each node, LLC is split into two partitions. One for LC workloads, and one for BE jobs. This allows the controller to minimize interference and thus increase average, cluster-wide, CPU utilization from $15\%$ to $60\%$.
@@
To improve the performance of accelerated ML workloads, (zhu_kelp_2019, 126) leverage non-uniform memory access (NUMA) node subdomains in their Kelp (zhu_kelp_2019, 126) runtime. Each system socket is split into two subdomains, one for high-priority ML workloads and one for low-priority tasks. This separation in combination with heuristic resource partitioning reduces task interference such that ML workload throughput is increased.
@@ 5.2.4. Workload Clustering Approaches
Themis (tang_themis_2023, 98) by (tang_themis_2023, 98) increases QoS and fairness of guest VMs by observing their memory bandwidth usage behavior. The VMs are classified into groups using the Euclidean K-Means algorithm, where the number of groups equal the number of available COS on the system. The controller then proactively partitions LLC and MB by following a heuristic algorithm. The classification and partitioning scheme is reevaluated over time by continuously observing VM bandwidth usage. The controller may also react to saturation of the memory bus by moving VMs to other NUMA nodes if available.
By observing a core’s LLC cache miss percentage and number of retired instructions, (gifford_dna_2021, 34) (gifford_dna_2021, 34) classify workloads into clusters depending on the program phase they are deemed to currently be in. These clusters are then allocated LLC ways and memory bandwidth [^5] Memory bandwidth is allocated using MemGuard (yun_memguard_2013, 118), a software-based approach. in order to maximize the performance of a system based on a heuristic algorithm. They call their approach DNA. A deadline-aware version that is able to take a real-time workload’s execution deadlines into consideration called DADNA is also presented. Their experimental evaluation of the controller only compares DNA and DADNA against (xu_holistic_2019, 112), another controller previously proposed by the same authors.
With the goal of increasing overall system performance, (xiang_emba_2019, 109) (xiang_emba_2019, 109) propose EMBA. EMBA is a controller that clusters workloads using a heuristic hierarchical clustering algorithm based on their usage of the memory bus. This classification is then used to heuristically and iteratively achieve a memory bandwidth throttle scheme that improves overall performance.
With system fairness as their goal, (selfa_application_2017, 92) (selfa_application_2017, 92) propose a set of LLC allocation policies that are based on workload clustering using the K-Means algorithm (hartigan_algorithm_1979, 38). The classification is done based on the number of stalled CPU cycles due to L2 cache misses. Then, the LLC is partitioned accordingly which results in an improvement of overall system fairness.
By clustering workloads using a support vector machine (SVM) based machine learning model, (qiu_machine-learning-based_2022, 87) are able to classify workload behavior using only system metrics. This is the first phase of their Clustering-and-Allocation (C&A) (qiu_machine-learning-based_2022, 87) LLC partitioning scheme. Using the aforementioned workload classification as input, a Bayesian optimization-based approach is used to quickly and efficiently reach a near-optimum partitioning scheme. In their own evaluation, C&A is found to increase throughput by up to $26.2\%$ in a best-case scenario.
@@ 5.2.5. Feature Abstractions
Due to the nature of hardware features such as Intel CAT, they are limited to bare-metal deployments with direct access to hardware. In an attempt to bring LLC allocation to virtualized environments, (xu_vcat_2017, 113) (xu_vcat_2017, 113) propose vCAT. vCAT is a system that exposes virtual cache partitioning to VMs, which is later mapped to physical LLC partitions on the host. By the extension of the functionality of hypervisors, guests are able to give hints regarding their workloads’ requirements, while the host remains in control.
@@ 5.2.6. Summary
Hardware QoS enforcement features may be used to great effect and tuned in many ways even when only system metrics are observed. Throughput may be increased by $26.2\%$ (qiu_machine-learning-based_2022, 87) or $18.5\%$ (chen_drlpart_2021, 16), performance increased by $7.1\%$ (navarro-torres_balancer_2023, 67), or fairness improved by $44\%$ (pons_cache-poll_2022, 78) compared to an unmanaged system.
Many publications classify workloads with regard to their utilization and sensitivity to a shortage of shared system resources. This categorization is then often used to dictate how resources are partitioned among workloads. Other papers try to predict how changes in resource allocation would impact the performance of workloads through machine learning and mathematical approaches. Conversely, reactive controllers have been proposed that continuously profile the workloads and reactively adjust resource allocations as requirements change over time. Additionally, publications focused on security (liu_catalyst_2016, 60), compiler inserted cache behavior hints (chatterjee_com-cas_2023, 12), auctioning-based partitioning (funaro_ginseng_2016, 31), and LLC management on the data-center scale (shahrad_provisioning_2021, 93), and more have been discussed.
@@ 5.3. Application Aware Approaches
To better understand how shared resource allocation affects the real-world performance of workloads, controllers may monitor key performance indicators (KPIs) of applications, see Figure 11. Such metrics may include request throughput, average latencies, tail latencies, and more. Every workload’s KPIs are different and manual configuration of metric probes is therefore often required. Resource management controllers need to be aware of how to measure an application’s metrics, as well as their target values and the importance of one metric compared to another. Many workloads have service-level objectives (SLOs) to adhere to which allow application-aware resource managers to minimize the resource availability of such workloads while ensuring they still operate within their targets. Thus, freeing more resources to maximize the performance of other co-located workloads.
Figure 11. Illustration of application-aware approaches and how they rely on application KPIs, sometimes in addition to system metrics. KPIs may include request throughput, average latencies, tail latencies, etc.:
We group application-aware approaches into four categories: Proactive, reactive, combined, and performance slack management approaches. See Table 5.
@@
(jeatsa_casy_2022, 51) Cloud Performance ML Resource Management Controller for FaaS
(stewart_performance_2022, 97) Generic Performance ML Short-term LLC Allocation Controller
(noll_accelerating_2018, 71) DBMS Performance LLC Management for DBMS
@@
(nikas_dicer_2019, 70) Generic Utilization Heuristic LLC Partitioning Scheme
(chintapalli_restrain_2022, 18) VNFs Performance Heuristic VNF Resource Management
(duan_improving_2021, 24) Generic Performance ML Classification and Heuristic Scheduling
(duan_improving_2022, 25) Generic Performance ML Classification and Heuristic Scheduling
(patel_clite_2020, 76) Cloud Performance Bayesian Optimization Based Algorithm
(pons_effect_2022, 81) Generic Performance Analysis of App Sensitivity to Shared Resources
(gureya_generalizing_2021, 36) Cloud Performance Heuristic Multi-Socket Memory Management
(papadakis_improving_2017, 73) Generic Performance Heuristic LLC Partitioning Scheme
(qin_locpart_2017, 83) Generic Performance Heuristic LLC Partitioning Scheme
@@
(qiu_slo_2021, 85) Generic Performance Profiling & Feedback Based Management
(nejat_coordinated_2020, 69) Generic Energy Efficiency Modeling of Workload Behavior
(zhao_rhythm_2020, 123) Cloud Utilization Sub-service LLC Management Scheme
(tootoonchian_resq_2018, 102) VNFs Utilization Greedy Profiling Based LLC Allocation
(metsch_intent-driven_2023, 64) Cloud Performance ML and Intent-driven Resource Management
(li_rldrm_2020, 57) VNFs Utilization Deep Reinforcement Learning LLC Management
(zhang_libra_2021, 122) Generic Performance Heuristic Memory Bandwidth Management
(chintapalli_ravin_2023, 19) VNFs Utilization Heuristic VNF and Shared Resource Allocation
(xu_holistic_2019, 112) RT Performance Heuristic LLC and BW Management Algorithm
@@
(chen_parties_2019, 17) Cloud Utilization Heuristic Slack Management Algorithm
(zhu_dirigent_2016, 125) Generic Utilization Execution Time Prediction Algorithm
(lo_heracles_2015, 61) Cloud Performance Heuristic Slack Management Algorithm
(nejat_cooperative_2022, 68) Generic Energy Efficiency Heuristic Slack Management Algorithm
(pang_adaptive_2021, 72) Cloud Energy Efficiency Multi-level Resource Management Scheme
(chen_avalon_2019, 14) Generic Utilization Runtime Prediction and LLC Management
@@
As one of the more recognized and important works in the category, PARTIES (chen_parties_2019, 17) is a QoS-aware resource management controller that allocates and tunes LLC, CPU cores, CPU frequency, and disk I/O to improve overall performance. It works with a single LC workload co-located with multiple BE jobs that initially are allocated equal amounts of resources. The controller then observes each application’s tail latency to determine its performance slack in relation to a predefined QoS target. The applications with small or negative slack is selected for upsizing, where the controller iteratively tries to increase allocations one resource at a time until the slack is increased. Should an increase in an allocation not lead to increased slack, the change is reverted. This is similarly repeated for applications with large slack, but in the opposite direction through downsizing. This process is repeated while applications fulfill predefined criteria for either upsizing or downsizing and is illustrated in Figure 12. In the event that the controller is unable to find a configuration that satisfies all applications’ SLOs, it will select an appropriate application for node migration (chen_parties_2019, 17).
Figure 12. Illustration of the iterative resource adjustment scheme of PARTIES by (chen_parties_2019, 17) (chen_parties_2019, 17).: